OpenStudy (anonymous):

Consider the differential equation dy/dx = cos(xpi)(y-1)^2 a) There is a horizontal line with equation y = c that satisfies this differential equation. Find the value of c.

OpenStudy (anonymous):

\[dy/dx= cos(\pi x)(y-1)^2\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{dy}{(y-1)^2}=cos(\pi x)dx\] integrate both sides \[\frac{1}{1-y}= \frac{1}{\pi}sin(\pi x) +C\]

OpenStudy (anonymous):

I don't understand the y = c part satisfies the equation. What does that mean?

OpenStudy (mr.math):

Can you solve the equation imran wrote for y?

OpenStudy (anonymous):

I just realized that this might be easier than what I was doing y=c dy/dx=0 0 = cos(x pi)(c-1)^2

OpenStudy (anonymous):

and solving for c , we get c=1 or c=-1

OpenStudy (mr.math):

\[y=1-\frac{\pi}{\pi{c}+\sin(\pi{x})}.\]

OpenStudy (mr.math):

imran's approach could be easier.

OpenStudy (anonymous):

\[\text{Cos}[\pi x]-2 c \text{Cos}[\pi x]+c^2 \text{Cos}[\pi x]\text{==}0\] \[1-2 c +c^2 \text{==}0\] (c+1)(c-1)=0 c=1,-1

OpenStudy (mr.math):

Never mind! I was stupid :D Go with imran.

OpenStudy (anonymous):

Ah, I don't really get why we differentiated y = c. I understand plugging it in though.

OpenStudy (anonymous):

because we wanted y'

OpenStudy (anonymous):

So when they say that y = c satisfies the equation. It means that y' = the differential equation?

OpenStudy (anonymous):

if y=constant y' must be zero